# Resilient Topology-Aware Coordination for Dynamic 3D UAV Networks under Node Failure

### v1

@@ 6.3. Performance Metrics
To quantify the system’s performance and its adaptive behavior under disruptions, we evaluate the following four key metrics, leveraging the formal definitions provided in Section 4:
@@
  • Network Resilience ($C_{\text{cov}}$ & $\mathcal{J}_{\text{R}}$): We evaluate the post-failure recovery of the Coverage Ratio and Jain’s Fairness Index as defined in Sections 4.1.3 and 4.1.4, respectively. These serve as the primary indicators of the swarm’s “survivability” and its autonomous ability to re-equilibrate service distribution after a loss of hardware resources.
@@
  • Topological Stability ($C_{\text{HO}}$): This represents the total handoff count per episode. As defined in Section 4.1.5, this metric is critical for assessing whether the controller maintains service through stable spatial positioning or via erratic user-UAV re-associations that would trigger signaling storms in real-world 6G deployments.
@@
  • Operational Efficiency ($E_{\text{eff}}$): Defined in Section 4.1.2, this global energy efficiency metric reflects the ratio of system throughput to the total power consumption of the active UAV set $\mathcal{V}(t)$. It evaluates the sustainability of the swarm’s movement and service strategy.
@@
  • System Capacity ($R_{\text{sum}}$): The aggregated system throughput as defined in Section 4.1.1. This metric reflects the swarm’s capability to mitigate co-channel interference and manage the increased load-per-agent following a node failure in dense environments.
6.4. Training Convergence and QoS Optimization
The training trajectories reflecting learning efficacy and Quality of Service (QoS) are illustrated in Figure 3. In this context, the Average Episode Reward and Coverage Ratio ($C_{\text{cov}}$) serve as primary indicators for evaluating the models’ progress across heterogeneous environments.
6.4.1. Learning Efficacy and Convergence Rate
As depicted in the top row of Figure 3, TAG-MAPPO exhibits superior convergence speed and asymptotic performance compared to both G-MAPPO (MLP baseline) and QMIX. Specifically, in the Crowded Urban scenario shown in Figure 3(a), TAG-MAPPO achieves a steady-state reward of approximately 2.2. This result surpasses the MLP baseline by 10% and significantly outperforms the geometrical K-Means approach, which reaches only 1.6.
This performance gap is most pronounced in interference-limited regimes, as the topology-aware critic successfully deciphers the complex inter-agent coordination requirements needed to minimize overlap while maximizing cluster coverage. In the Suburban and Rural cases, represented by Figure 3(b) and Figure 3(c) respectively, TAG-MAPPO maintains its lead despite the increased environmental dispersion. Notably, the framework demonstrates a significant reduction in the variance of the learning process, indicating higher training stability.
6.4.2. QoS Performance and Spatial Coverage
The evolution of $C_{\text{cov}}$ in the bottom row of Figure 3 provides a physical interpretation of these reward gains. For the Suburban deployment illustrated in Figure 3(e), TAG-MAPPO stabilizes at a coverage ratio of 0.55 within the first 1500 episodes while maintaining a tighter confidence interval than its competitors. This suggests that graph-based feature aggregation is inherently less sensitive to the stochasticity of user mobility compared to traditional neural architectures.
The robustness of the framework is further evidenced in Rural Areas as shown in Figure 3(f). Although independent Gauss-Markov mobility leads to higher reward fluctuations, our framework effectively prevents coverage collapse. The integration of Residual Ego-State Fusion ensures that each agent maintains sufficient self-awareness to track isolated user clusters even when neighbor information is sparse. These findings confirm that the proposed framework provides robust and scalable intelligence for 3D AGINs across diverse geographical profiles.
@@ 6.4.2. QoS Performance and Spatial Coverage
Fig 3. Training convergence analysis across heterogeneous scenarios. The top row, including Figs. 3(a)3(c), illustrates the Average Episode Reward, while the bottom row, Figs. 3(d)3(f), depicts the evolution of the Coverage Ratio ($C_{\text{cov}}$). Results are averaged over 5 independent runs, with shaded areas representing the 95% confidence interval, demonstrating the framework’s stability under stochastic mobility.
6.5. Stability and Resource Efficiency Analysis
The operational sustainability and signaling stability of the TAG-MAPPO framework are evaluated through the Total Handoffs ($C_{\text{HO}}$) and Energy Efficiency ($E_{\text{eff}}$), as illustrated in Figure 4.
6.5.1. Topological Stability and Handoff Suppression
The top row of Figure 4 demonstrates the models’ ability to maintain service continuity under the handoff penalty $\lambda_{\text{HO}}$. As evidenced by the results for the Crowded Urban and Suburban environments in Figure 4(a) and Figure 4(b), TAG-MAPPO consistently maintains the lowest handoff frequency, stabilizing below 50 handoffs per slot. In contrast, the G-MAPPO (MLP baseline) exhibits significant fluctuations and a higher handoff baseline. This performance gap is especially evident in the Suburban scenario, where G-MAPPO frequently exceeds 100 handoffs.
A critical insight is observed in the Rural Area shown in Figure 4(c). While QMIX and G-MAPPO suffer from a handoff explosion after episode 1500, TAG-MAPPO remains remarkably suppressed. Notably, G-MAPPO peaks at nearly 400 handoffs due to erratic re-associations in sparse topologies. These observations prove that the topology-aware critic effectively discourages ping-pong handoffs by enabling UAVs to learn long-term spatial coordination. This strategy prioritizes stable link associations over short-sighted coverage gains, which is essential for reducing signaling overhead.
6.5.2. Resource Utilization and Energy Efficiency
The bottom row of Figure 4 evaluates the system’s throughput-per-Watt performance. In the Crowded Urban scenario represented by Figure 4(d), TAG-MAPPO stabilizes at approximately 5 Mbps/W, maintaining a clear lead over the K-Means geometrical baseline of 3.4 Mbps/W. The learning efficiency is further highlighted in Figure 4(e), where the energy efficiency of TAG-MAPPO converges much faster than QMIX, which requires over 2000 episodes to approach a similar performance level.
The integration of graph-based feature aggregation allows agents to identify energy-optimal hovering positions that minimize propulsion power while maximizing communication utility. As evidenced across all geographical profiles from Figure 4(d) to Figure 4(f), TAG-MAPPO not only achieves superior efficiency but also maintains a much tighter confidence interval than MLP-based approaches. This consistent resource efficiency is vital for the Internet of UAVs. It ensures battery longevity and minimizes control-plane overhead for mission-critical 6G services.
@@ 6.5.2. Resource Utilization and Energy Efficiency
Fig 4. Stability and Energy Efficiency Analysis: The top row, including Figs. 4(a)4(c), illustrates the Total Handoffs ($C_{\text{HO}}$) across different scenarios, while the bottom row, Figs. 4(d)4(f), depicts the Energy Efficiency ($E_{\text{eff}}$). Results are averaged over 5 independent runs, with shaded areas representing the 95% confidence interval, demonstrating the framework’s superior stability and resource efficiency under stochastic mobility.
6.6. System Resilience under Catastrophic Node Failure
To evaluate the survivability of the proposed framework, we analyze the system performance following a random node failure at $t=100$. A fundamental advantage of TAG-MAPPO is its inherent capability to handle dynamic changes in the agent population. Unlike conventional MARL architectures such as QMIX, which rely on a centralized mixing network with fixed input dimensions, our framework leverages the permutation-invariant nature of graph-based feature aggregation. In a QMIX-based system, the loss of a node at $t=100$ leads to a structural mismatch in the mixing layer, making it unsuitable for resilience tests without manual reconfiguration. In contrast, TAG-MAPPO processes the network as a graph of relational entities, allowing the surviving UAVs to maintain stable coordination even when a vertex is abruptly removed. This flexibility is visualized in Figure 5, where we focus on the comparative analysis between TAG-MAPPO and the G-MAPPO baseline.
6.6.1. Dynamic Recovery and Coverage Self-healing
As illustrated in the top row of Figure 5, TAG-MAPPO exhibits a distinct V-shaped recovery trajectory across all scenarios immediately after the node failure. While the Coverage Ratio experiences an abrupt decline at $t=100$, the surviving agents autonomously reconfigure their spatial distribution to compensate for the service gap. For the Crowded Urban scenario shown in Figure 5(a), TAG-MAPPO restores its coverage level to near pre-failure states significantly faster than the baseline.
The most prominent advantage is observed in the Suburban and Rural Area cases illustrated in Figure 5(b) and Figure 5(c). In these environments, the baseline coverage continues to fluctuate or degrade after the initial shock, whereas TAG-MAPPO stabilizes the network rapidly. This resilience is attributed to the Random Observation Shuffling (ROS) mechanism. This training technique ensures the policy remains robust against unexpected changes in agent count by preventing the remaining agents from over-relying on a fixed set of neighbors.
6.6.2. Structural Reorganization and Fairness Improvement
Beyond mere coverage restoration, the bottom row of Figure 5 highlights the evolution of Jain’s Fairness Index (JFI), providing deeper insights into how resource distribution is maintained under stress. A remarkable phenomenon is observed in the Crowded Urban scenario in Figure 5(d). Following the node failure, the JFI of TAG-MAPPO not only recovers but actually surpasses its pre-failure level. This suggests that the topology-aware policy successfully identifies and resolves the inter-agent service overlaps that existed in the four-UAV configuration. By re-optimizing the spatial deployment of the three remaining agents, the framework achieves a more balanced resource distribution across the user plane.
In the Rural Area shown in Figure 5(f), the baseline JFI collapses towards 0.25 after the node failure, indicating that the MLP-based agents fail to reorganize their coverage in sparse environments. In contrast, TAG-MAPPO maintains a significantly higher JFI with a tighter confidence interval. These results confirm that the graph-based reasoning enables UAVs to maintain service survivability and fairness even in hostile operational conditions. The integration of Residual Ego-state Fusion ensures that each agent retains sufficient self-awareness to track user clusters independently, which is a critical requirement for 6G mission-critical reliability during hardware malfunctions.
@@ 6.6.2. Structural Reorganization and Fairness Improvement
Fig 5. System Resilience Analysis: The top row, including Figs. 5(a)5(c), illustrates the Coverage Ratio ($C_{\text{cov}}$) following a random node failure at $t=100$, while the bottom row, Figs. 5(d)5(f), depicts the Jain’s Fairness Index (JFI) over the same period. Results are averaged over 5 runs with shaded areas representing the standard deviation.
7. Conclusion
@@
In this paper, we presented TAG-MAPPO, a topology-aware reinforcement learning framework designed to optimize Quality of Service (QoS) and operational stability in 3D Aerial-Ground Integrated Networks (AGINs). By integrating graph-based feature aggregation with a residual ego-state fusion mechanism, the proposed framework successfully deciphers complex spatial dependencies among UAV agents in highly dynamic environments. Our approach addresses the fundamental challenges of signaling overhead and resource inefficiency that often hinder the deployment of autonomous aerial base stations.
Experimental results across three heterogeneous geographical profiles demonstrate that TAG-MAPPO significantly outperforms conventional MLP-based MAPPO and QMIX baselines. Notably, the framework achieves a superior balance between coverage expansion and handoff suppression; it reduces redundant signaling by up to 50% in sparse rural topologies while maintaining high energy efficiency. Furthermore, the system resilience tests highlight the unique self-healing capability of TAG-MAPPO. Following a catastrophic node failure, the network autonomously reconfigures its topology to restore over 90% of the pre-failure service coverage within 15 time steps. In dense urban scenarios, the framework even achieves a higher Jain’s Fairness Index post-failure compared to its original four-UAV configuration by resolving service overlaps. These findings confirm that incorporating topological intelligence is essential for achieving robust and scalable coordination in 6G mission-critical missions.
For future work, we aim to extend this framework to multi-tier AGIN architectures involving satellite-UAV coordination. Additionally, investigating the integration of secure communication protocols within the topology-aware learning process would be a valuable direction to ensure the integrity of the Internet of UAVs against adversarial interference.
Acknowledgements.
This work was supported by the National Science and Technology Council (NSTC), Taiwan, under Grant No. NSTC 114-2221-E-194-062-, and was also partially supported by the Advanced Institute of Manufacturing with High-tech Innovations (AIM-HI) from the Featured Areas Research Center Program within the framework of the Higher Education Sprout Project by the Ministry of Education (MOE) in Taiwan.

### v2

@@ VI-C Performance Metrics
To quantify the system’s performance and its adaptive behavior under disruptions, we evaluate the following four key metrics, leveraging the formal definitions provided in Section IV:
@@
  • Network Resilience ($C_{\text{cov}}$ & $\mathcal{J}_{\text{R}}$): We evaluate the post-failure recovery of the Coverage Ratio and Jain’s Fairness Index as defined in Sections IV-A3 and IV-A4, respectively. These serve as the primary indicators of the swarm’s “survivability” and its autonomous ability to re-equilibrate service distribution after a loss of hardware resources.
@@
  • Topological Stability ($C_{\text{ho}}$): This represents the total handoff count per episode. As defined in Section IV-A5, this metric is critical for assessing whether the controller maintains service through stable spatial positioning or via erratic user-UAV re-associations that would trigger signaling storms in real-world 6G deployments.
@@
  • Operational Efficiency ($E_{\text{ee}}$): Defined in Section IV-A2, this global energy efficiency metric reflects the ratio of system throughput to the total power consumption of the active UAV set $\mathcal{V}(t)$. It evaluates the sustainability of the swarm’s movement and service strategy.
@@
  • System Capacity ($R_{\text{sum}}$): The aggregated system throughput as defined in Section IV-A1. This metric reflects the swarm’s capability to mitigate co-channel interference and manage the increased load-per-agent following a node failure in dense environments.
@@
Figure 3: Training convergence analysis across heterogeneous scenarios. The top row, including Figs. 3(a)3(c), illustrates the Average Episode Reward, while the bottom row, Figs. 3(d)3(f), depicts the evolution of the Coverage Ratio ($C_{\text{cov}}$). Results are averaged over 5 independent runs, with shaded areas representing the 95% confidence interval, demonstrating the framework’s stability under stochastic mobility.
VI-D Training Convergence and QoS Optimization
The training trajectories, reflecting learning efficacy and multi-objective optimization across three distinct mobility phases, are illustrated in Fig. 3. To evaluate the framework’s robustness, the 5,000-episode training process is divided into: 1) a standard mobility phase ($0\leq ep<1600$), 2) a dynamic merging phase ($1600\leq ep<3300$), and 3) a high-speed evacuation phase ($3300\leq ep\leq 5000$).
VI-D1 Learning Efficacy and Convergence Rate
As depicted in the top row of Fig. 3, the Average Episode Reward exhibits distinct step-like transitions corresponding to the shifts in user mobility patterns. In the Crowded Urban and Suburban scenarios shown in Fig. 3(a) and Fig. 3(b), the reward abruptly oscillates at $ep=1600$ as users converge toward the map center. TAG-MAPPO demonstrates superior adaptability, rapidly recalibrating its policy to exploit the increased user density for higher signal-to-interference-plus-noise ratio (SINR).
At $ep=3300$, the transition to the high-speed evacuation mode ($15\text{ m/s}$) triggers a reward drop across all models. However, TAG-MAPPO exhibits a faster recovery slope compared to the G-MAPPO (MLP baseline) as illustrated in Fig. 3(c). This advantage stems from the topology-aware attention mechanism, which allows agents to maintain spatial correlation even when user clusters dissipate rapidly. Notably, the QMIX baseline fails to converge in these highly dynamic regimes, underscoring the necessity of policy-based coordination in 3D aerial networks.
VI-D2 Resilience and Coverage Optimization
The evolution of the Coverage Ratio $C_{\text{cov}}$ in the bottom row of Fig. 3 provides a physical interpretation of the system’s resilience. Despite the inclusion of random node failures during training, TAG-MAPPO maintains a stable $C_{\text{cov}}$ around 0.350.350.450.45 in urban environments. A significant observation is the reduced variance, represented by the shaded area, of TAG-MAPPO compared to the MLP baseline during the high-speed phase where $ep>3300$. This indicates that graph-based feature aggregation is inherently more robust to the distribution shifts caused by rapid topological changes.
In Rural Areas, as illustrated in Fig. 3(c) and Fig. 3(f), where user sparsity poses a greater challenge for collaborative tracking, TAG-MAPPO maintains a consistent lead. While G-MAPPO exhibits significant performance collapse in the final stage, TAG-MAPPO effectively prevents coverage voids through its Residual Ego-State Fusion, ensuring that agents retain local mission awareness even when neighbor information is sparse. These findings confirm that the proposed framework provides the necessary intelligence for 3D AGINs to maintain Quality of Service (QoS) amidst extreme mobility and unpredictable hardware failures.
@@ VI-D2 Resilience and Coverage Optimization
Figure 4: Stability and Energy Efficiency Analysis: The top row, including Figs. 4(a)4(c), illustrates the Total Handoffs ($C_{\text{ho}}$) across different scenarios, while the bottom row, Figs. 4(d)4(f), depicts the Energy Efficiency ($E_{\text{ee}}$). Results are averaged over 5 independent runs, with shaded areas representing the 95% confidence interval, demonstrating the framework’s superior stability and resource efficiency under stochastic mobility.
VI-E Stability and Resource Efficiency Analysis
The operational sustainability and signaling stability of the TAG-MAPPO framework are evaluated through the Total Handoffs $C_{\text{ho}}$ and Energy Efficiency $E_{\text{ee}}$ as illustrated in Fig. 4.
VI-E1 Topological Stability and Handoff Suppression
The top row of Fig. 4 illustrates the framework’s capability to suppress excessive signaling overhead amidst environmental transitions. As evidenced in the Crowded Urban and Suburban scenarios shown in Fig. 4(a) and Fig. 4(b), TAG-MAPPO maintains a remarkably low and stable handoff frequency, consistently staying below 50 events per slot.
A critical observation occurs at the $ep=3300$ transition to high-speed evacuation where the velocity is $15\text{ m/s}$. At this junction, the G-MAPPO baseline suffers from a handoff explosion that peaks near 200 in Rural Areas, as depicted in Fig. 4(c). In contrast, TAG-MAPPO remains suppressed throughout the phase. This disparity suggests that the topology-aware attention mechanism allows agents to distinguish between necessary proactive re-associations and erratic ping-pong handoffs caused by rapid user dispersion. Furthermore, the QMIX and MLP baselines exhibit high variance, represented by the shaded areas in Fig. 4, indicating that traditional architectures struggle to maintain connection stability when the underlying graph topology undergoes rapid expansion.
VI-E2 Resource Utilization and Energy Efficiency
The bottom row of Fig. 4 evaluates the system’s throughput-per-Watt performance. In the Crowded Urban scenario shown in Fig. 4(d), TAG-MAPPO achieves a steady-state efficiency of approximately 5 Mbps/W, which significantly outperforms the K-Means geometrical baseline of 3.4 Mbps/W. This gain is maintained even during the merging phase where $1600\leq ep<3300$, a period when high user density typically complicates interference management.
The robustness of the proposed framework is further highlighted in the high-speed regime for $ep>3300$. As shown across all scenarios in Fig. 4(d) through Fig. 4(f), TAG-MAPPO maintains a tighter confidence interval and higher asymptotic $E_{\text{ee}}$ compared to its competitors. By integrating graph-based features, agents can identify energy-optimal hovering positions that minimize high-speed propulsion power $P_{k}^{\text{fly}}$ while maximizing communication utility. This consistency in resource efficiency is vital for mission-critical 6G services, as it ensures both battery longevity and control-plane stability in volatile aerial environments.
@@ VI-E2 Resource Utilization and Energy Efficiency
Figure 5: System Resilience Analysis: The top row, including Figs. 5(a)5(c), illustrates the Coverage Ratio ($C_{\text{cov}}$) following a random node failure at $t=100$, while the bottom row, Figs. 5(d)5(f), depicts the Jain’s Fairness Index (JFI) over the same period. Results are averaged over 5 runs with shaded areas representing the standard deviation.
VI-F System Resilience under Catastrophic Node Failure
To evaluate the survivability of the proposed framework, we analyze the system performance following a random node failure at $t=100$. A fundamental advantage of TAG-MAPPO is its inherent capability to handle dynamic changes in the agent population. Unlike conventional MARL architectures such as QMIX, which rely on a centralized mixing network with fixed input dimensions, our framework leverages the permutation-invariant nature of graph-based feature aggregation.
In a QMIX-based system, the loss of a node at $t=100$ leads to a structural mismatch in the mixing layer, which makes it unsuitable for resilience tests without manual reconfiguration. In contrast, TAG-MAPPO processes the network as a graph of relational entities, allowing the surviving UAVs to maintain stable coordination even when a vertex is abruptly removed. This flexibility and the subsequent recovery process are visualized in Fig. 5.
VI-F1 Dynamic Recovery and Coverage Self-healing
As illustrated in the top row of Fig. 5, TAG-MAPPO exhibits an immediate and robust response to the catastrophic event at $t=100$. In the Crowded Urban and Suburban scenarios shown in Fig. 5(a) and Fig. 5(b), the Coverage Ratio $C_{\text{cov}}$ does not suffer from a persistent collapse. Instead, it maintains a stable level that significantly outperforms the G-MAPPO baseline.
This superior resilience is achieved through the integration of the Random Observation Shuffling (ROS) mechanism during the training phase. By stochastically permuting the sequence of neighbor observations, ROS prevents the agents from developing a coordinate-index dependency where the policy implicitly relies on the fixed ordering of neighbor information. Consequently, when a node failure occurs, the surviving agents do not perceive the missing data as a broken input but rather as a valid topological variation.
This structural generalization, reinforced by exposure to stochastic failures during training, allows the swarm to internalize a robust coordination logic that adapts to unpredictable topological ruptures. Upon the failure of a node, the remaining agents immediately re-evaluate the transformed topology through the attention mechanism to fill the service gap, ensuring that the aggregate $C_{\text{cov}}$ remains competitive despite the 25 percent reduction in available hardware resources.
VI-F2 Structural Reorganization and Performance Gain
Beyond mere coverage restoration, the bottom row of Fig. 5 highlights a remarkable phenomenon in Jain’s Fairness Index $\mathcal{J}_{\text{r}}$. Following the node failure in the Crowded Urban scenario depicted in Fig. 5(d), the $\mathcal{J}_{\text{r}}$ of TAG-MAPPO actually surpasses its pre-failure level, rising from 0.35 to 0.6.
This counter-intuitive performance gain suggests that the proposed framework achieves a more efficient spatial equilibrium after the failure. In high-density environments, the removal of a redundant or interfering node allows the surviving agents to eliminate service overlaps and coordination conflicts. The ROS mechanism plays a pivotal role here by ensuring that the feature aggregation logic treats the neighboring swarm as a flexible set rather than a fixed vector. By autonomously re-optimizing their positions, the remaining swarm mitigates co-channel interference that was sub-optimally handled in the original four-agent configuration. This achievement of a more efficient spatial equilibrium is particularly significant as it demonstrates the model’s ability to identify and resolve service redundancies in real time.
In contrast, the results for the Rural Area shown in Fig. 5(f) show that the MLP baseline fails to reorganize, with its $\mathcal{J}_{\text{r}}$ collapsing towards 0.25. This failure indicates that traditional architectures struggle to maintain connection stability when the underlying graph topology undergoes rapid contraction. TAG-MAPPO, however, maintains a significantly higher fairness index with a tighter confidence interval. These results confirm that graph-based reasoning enables UAVs to maintain service survivability even in hostile operational conditions. The integration of residual ego-state fusion ensures that each agent retains sufficient self-awareness to track user clusters independently, satisfying the critical requirements for 6G mission-critical reliability during hardware malfunctions.
VII Conclusion
@@
This paper presented TAG-MAPPO, a topology-aware reinforcement learning framework designed to optimize service quality and operational stability in 3D Aerial-Ground Integrated Networks. By integrating graph-based feature aggregation with a residual ego-state fusion mechanism, the framework successfully deciphers complex spatial dependencies among UAV agents in dynamic environments. Our approach addresses fundamental challenges of signaling overhead and resource inefficiency that hinder autonomous aerial base station deployment in 6G scenarios. Experimental results across three geographical profiles demonstrate that TAG-MAPPO outperforms conventional MLP-based MAPPO and QMIX baselines. By implementing the Random Observation Shuffling (ROS) mechanism, the framework reduces redundant signaling by up to 50 percent in sparse rural topologies while maintaining superior energy efficiency and stability during high-speed mobility phases.
The system resilience tests highlight the structural generalization capability of TAG-MAPPO, where the surviving swarm autonomously reconfigures its topology to restore coverage immediately following catastrophic node failures. In dense urban scenarios, the framework achieves a higher fairness index post-failure compared to its original four-UAV configuration, illustrating the attention mechanism’s ability to resolve service overlaps once interference sources are removed. These findings confirm that topological intelligence is essential for robust and scalable coordination in 6G mission-critical missions. Future work will extend this framework to multi-tier architectures involving satellite-UAV coordination. Additionally, integrating secure communication protocols within the topology-aware learning process remains a valuable direction to ensure the integrity of the Internet of UAVs against adversarial interference.